### v5

# Enhancing User Interest based on Stream Clustering and Memory Networks in Large-Scale Recommender Systems

Abstract.
Recommender Systems (RSs) provide personalized recommendation service based on user interest, which are widely used in various platforms. However, there are lots of users with sparse interest due to lacking consumption behaviors, which leads to poor recommendation results for them. This problem is widespread in large-scale RSs and is particularly difficult to address. To solve this challenging problem, we propose an innovative solution called User Interest Enhancement (UIE). UIE enhances user interest including user profile and user history behavior sequences by leveraging the enhancement vectors and personalized enhancement vectors generated based on dynamic streaming clustering of similar users and items from multiple perspectives, which are stored and updated in memory networks. UIE not only remarkably improves model performance for users with sparse interest, but also delivers notable gains for other users. As an end-to-end solution, UIE is easy to implement on top of existing ranking models. Furthermore, we extend our approach to long-tail items using similar methods, which also yields excellent improvements.
We conduct extensive offline and online experiments in a large-scale industrial RS. The results demonstrate that our model substantially outperforms other existing approaches, especially for users with sparse interest. UIE has been deployed in several large-scale RSs at Tencent since 2022, which was made public on 21 May 2024. In addition, UIE-based methods have also been successfully applied in candidate generation, pre-ranking, and context-DNN stages. Multiple teams have developed solutions based on UIE, focusing primarily on updating clustering algorithms and attention mechanisms. As far as we know, UIE has been deployed by many companies. The thoughts of UIE, dynamic streaming clustering and similarity enhancement, have inspired subsequent relevant works.
Recommender Systems, User Interest Enhancement, Stream Clustering and Memory Networks, Long-Tail Optimization, Multi-Task Learning
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; May 21, 2024; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’24), May 21, 2024, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/24/05doi: 10.1145/3690624.3709429ccs: Information systems Recommender systemsccs: Information systems Personalization
1. Introduction
Recommender Systems (RSs) (ref1, ; ref2, ) which provide personalized recommendation service based on user interest are widely used in various platforms such as short video platforms (ref3, ; ref7, ; ref14, ), video platforms (ref4, ; ref5, ), E-commerce platforms (ref6, ; ref8, ; ref9, ; ref10, ; ref11, ) and social networks (ref12, ; ref13, ), serving billions of users. In RSs, ranking typically uses a Multi-Task Learning model (MTL) (ref4, ; ref8, ; ref16, ; ref17, ; ref18, ; ref19, ; ref20, ; ref21, ) and lots of features to finely predict the scores of multiple user behaviors such as valid click, watching time, fast slide, like and sharing for thousands of candidates. The accuracy of the scores outputted by MTL is crucial for RSs (ref4, ). In RSs, user interest includes user profile and user history behavior sequences, as shown in Table 1 and Figure 1, which determines the upper bound of ranking model’s performance. However, lots of users only have sparse interest due to lacking consumption behaviors. For example, among the Daily Active Users (DAU) of our RS, more than 30% of them have this problem, whose average consumption behaviors are less than half of that of other users. As a result, this makes it challenging for ranking model to accurately estimate scores for these users, leading to poor recommendation results for them. It is widespread in large-scale industrial RSs. Improving the quality of recommendations for the users with sparse interest is very important, but also particularly difficult.
Although there has been considerable research on ranking in RSs (ref4, ; ref8, ; ref16, ; ref17, ; ref18, ; ref19, ; ref20, ; ref21, ), studies focusing on users with sparse interest are relatively limited. (ref22, ; ref23, ; ref24, ; ref25, ; ref26, ; ref27, ) are designed to address cold start problem for new users and new items, making them unsuitable for users with sparse interest who have some consumption behaviors but not many. In the context of RSs, the data augmentation of Contrastive Learning (CL) can easily produce a significant amount of meaningless data, which offers limited value. Furthermore, the effectiveness of Transfer Learning (TL) (ref29, ) is heavily influenced by the discrepancies between source and target data distributions, which has limited potential. Additionally, (ref3, ; ref31, ) only optimize model structure and embedding whose effect are also not significant due to lack of critical information of users with sparse interest.
To improve model performance for users with sparse interest, we propose a novel solution called User Interest Enhancement (UIE). UIE enhances user interest—including user profile and historical behavior sequences—by enhancement vectors and personalized enhancement vectors generated with the help of similar users and relevant items based on stream clustering and memory networks. UIE significantly improves the performance of ranking models for users with sparse interest by supplementing pivotal interest information. Moreover, UIE also proves highly valuable for other users, achieving substantial improvements. UIE is an end-to-end solution and can be easily implemented on top of existing ranking models. Similarly, highly skewed long-tail item distribution are common in RSs, which significantly hurts model performance on tail items. Therefore, we extend our solution and apply similar methods to tail items, achieving excellent improvements as well. To the best of our knowledge, this is the first work that proposes the thoughts of dynamic streaming clustering and similarity enhancement for enhancing user interest, which have inspired subsequent relevant works.
The major contributions of our work include:
• We first point out the problem that a large number of users only have sparse interest in large-scale RSs, which leads to poor recommendation performance for them. Moreover, a similar problem also exists for tail items.
• We propose UIE, an innovative method designed to enhance user interest, which includes three components for enriching user profile and user historical behavior sequences, respectively. By supplementing key interest information using enhancement vectors generated based on stream clustering and memory networks from different perspectives, model performance for users with sparse interest is significantly improved. UIE also boosts performance for other users. Furthermore, we also apply similar methods to tail items.
• We conduct offline evaluations to compare our model with other models on the same dataset, and the results demonstrate that our model outperforms the others significantly. Furthermore, we conduct online A/B tests in a large-scale RS serving hundreds of millions of users, and the results show that our model performs significantly better than the others.
• UIE has been fully deployed in our RS and multiple other large-scale RSs since 2022. As an algorithm applied in large-scale industrial RSs, we prefer simple and efficient solutions to balance performance and cost, which is also the fundamental principle guiding our algorithm design, rather than excessively pursuing complexity. Additionally, UIE-based methods have proven effective in the candidate generation, pre-ranking, and context-DNN stages. Multiple teams have developed solutions based on UIE, mainly updating the clustering and attention algorithms.
2. Data Analysis and Definition
This section gives the definition of users with sparse interest in our recommendation scenario through analyzing users’ profile and consumption behaviors. As previously mentioned, user interest contains user profile and user history behavior sequences. User profile includes basic attributes such as user id, age, gender, hometown, etc. and his/her preference generated based on this user’s consumption behaviors over a period of one month or several months, as shown in Table 1. For simplicity, in this paper, we take user valid consumption sequence as an example of various user history behavior sequences, as shown in Figure 1. A valid consumption is watching a video more than 10 seconds.
Table 1. An example of user profile.
Basic Attribute user id, age, gender, hometown, phone brand, …
User Preference On Category 1 sports:260, technology:120, military:109, politics:90, …
User Preference On Category 2 baskball:210, football:130, aerospace:82, weapon:80, …
User Preference On Tag NBA:190, CBA:170, AIGC:100, space shuttle:80, tank:60, cannon:50, warplane:30, …
Others
Figure 1. An example of user valid consumption sequence.
Figure 2. System overview of our ranking model with UIE, which includes User Profile Enhancement, User Consumption Behavior Enhancement and User Consumption Sequence Enhancement designed for enhancing user interest. To generate fixed-length user_vec and item_vec, an auxiliary user-item model is defined. Memory networks are used for storage, retrieval and updating of information. All the enhancement vectors outputted by these enhancement components are used as new features of ranking model to supply additional interest information.
User interest decides the performance upper bound of ranking model, which is critical for personalized recommendation. We randomly select 10 million users and conduct analysis of their user interest distribution. We observe significant difference in the number of user profile features and consumption behaviors among different users. For convenience, we define users with sparse interest as low-active users and define other users as high-active users, which account for 32% and 68% respectively. Because the distributions of user interest and user consumption behaviors are commercially sensitive, we provide their comparative values between high-active users and low-active users, as shown in Table 2. The average number of user profile features of low-active users is only 54% of that of high-active users and the average number of consumption behaviors of low-active users is merely 41% of that of high-active users.
Due to low-active users severely lacking crucial interest information, the model performance on them is very poor. In our recommendation scenario, the test Area Under Curve (AUC) metric [32] of valid consumption of low-active users is approximately 2% lower than that of high-active users. Similarly, highly skewed long-tail item distribution widely exists in RSs. In our RS, among the items exposed to users in a day, the top 10% items account for over 90% of total exposures, which seriously hurts model performance on tail items.
Table 2. The comparative values of the average number of user profile features and the average number of consumption behaviors between low-active users and high-active users.
 User Group User Profile Consumption Behaviors
High-active Users 1.0 1.0
Low-active Users 0.54 0.41
3. Method
3.1. System Overview of Ranking Model
We use the Progressive Layered Extraction (PLE) model (ref18, ) with excellent performance as our ranking model for multi-task predictions, which include valid consumption task, watching time task and other tasks. User interest includes user profile and user history behavior sequences, which determines the upper bound of ranking model performance. To achieve better performance, we propose User Interest Enhancement (UIE) that includes three components: User Profile Enhancement (UPE), User Consumption Behavior Enhancement (UCBE) and User Consumption Sequence Enhancement (UCSE), which are designed for enhancing user profile and user consumption behavior sequences respectively, as shown in Figure 2. The components of UIE generate enhancement vectors for a user to supplement additional interest information from different perspectives, which are used as new features of ranking model to improve its performance, especially on low-active users. It should be noted that the hyperparameters of UIE need to be selected carefully, which have a great impact on model performance. An example of hyperparameter selection is given in Section 4. Memory networks (ref33, ) are used for storage, retrieval and updating of information efficiently. In addition, we employ similar approaches to tail items, to mitigate the problem of insufficient exposure of them.
3.2. User Profile Enhancement (UPE)
User profile contains user basic attributes and user preference on category 1, category 2, tag and so on which are produced based on a user’s consumption behaviors over a period, as shown in Table 1. As previously mentioned, more than 30% of daily active users in our recommendation scenario only have sparse user profile because of lacking consumption behaviors. To alleviate this problem, we design a novel approach called UPE to infer missing profile features for users with the help of other similar users, as shown in Figure 2. The basic idea of UPE is to enhance user profile for users using the most similar centroids stored in memory network generated based on end-to-end stream clustering. It is not only particularly effective for low-active users, but also effective for high-active users, which are proved by the offline and online experiments in Section 4. UPE includes the following key steps.
Firstly, getting the vector of user profile. To generate fixed length vectors, no matter how many features are taken as input, a simple auxiliary model is defined including user tower and item tower. The user tower and the item tower of this auxiliary model take the features of user profile and the features of item as input separately. The features of a user’s profile are produced based on this user’s consumption behaviors over a period of one month or more months, so they don’t change quickly. Moreover, to save resource, the embeddings of the auxiliary model’s features are obtained from the main model. And to ensure the performance of the main model, the gradients from the auxiliary model are not back-propagated to the main model. The vector outputted by user tower is called user profile vector (user_vec).
Secondly, end-to-end stream clustering based on similarity between user profile vectors and the centroids stored in memory network. Before training, the centroids of ${N}$ clusters are randomly initialized which are ${d}$-dimensional vectors stored in user memory network. The similarity scores between a user profile vector and all centroids are calculated by matrix multiplication efficiently. Then the user profile vector $\upsilon$ is clustered to the most similar cluster. In practice, we apply distillation approach to update the corresponding cluster centroid $\mu_{i}$ for improving update efficiency using a small update rate $\rho$, as shown in Formula 1. In addition, to balance the influence of different users on clustering and further reduce computational cost, we perform sampling based on user’s active type and about 30% of all samples participate in clustering.
(1) $\mu_{i}^{n+1}\longleftarrow arg\min_{\mu_{i}}\rho*\sum_{j=1}^{d}\left(\upsilon_{j}-\mu_{ij}^{n}\right)^{2}\,.$
Thirdly, retrieval the most similar centroids based on similarity for a user profile vector. The most ${K_{1}}$ similar centroids of user profile vector are retrieved based on similarity in user memory network. To ensure positive correlations between these selected centroids and user profile vector, it is necessary to filter each selected centroid based on whether its inner product with user profile vector is greater than zero. If the dot product is less than zero, this centroid will be replaced by zero vector.
Finally, generating user profile enhancement vector based on the most similar centroids and personal vector. After obtaining the ${K_{1}}$ most similar vectors, the weighted vector $\upsilon_{weighted}$ is calculated using attention mechanism, as shown in Formula 2. Furthermore, to achieve better personalized enhancement, a short length vector for each user is predefined called personal vector, which is used to achieve personalized fine-tuning of weighted vector. The weighted vector and personal vector are concatenated and used as the input of the generalization layer, as shown in Figure 2.
(2) $\upsilon_{weighted}=\sum_{j=1}^{k_{1}}\frac{e^{(\mu_{j}\ *\ \upsilon)}*\mu_{j}}{\sum_{j=1}^{k_{1}}e^{(\mu_{j}\ *\ \upsilon)}}.$
The clustering algorithm used in UPE can be replaced with other clustering algorithms. Furthermore, when resources are sufficient, clustering can be performed separately for each ranking task and different media types (for example, in scenarios where images, text, and videos are recommended), or even tree-structured multi-level clustering can be adopted to further improve performance.
3.3. User Consumption Behavior Enhancement (UCBE)
Considering that a user’s top preferences dominate his/her user profile and don’t change rapidly, as shown in Table 1, so the most similar centroids selected by the user’s profile vector will be determined by his/her top preferences and remain fixed over a period. To further enhance a user’s profile, we construct a personalized enhancement vector for each user from the perspective of user consumption behaviors, as a supplement to UPE, which is called UCBE, as shown in Figure 2. The key idea of UCBE is to select the most relevant centroids in user memory network using a user’s consumption behaviors which considers both the user’s top preferences and tail preferences to construct a personalized enhancement vector. The key steps of UCBE are as follows.
Firstly, getting the vector of item features. We select the sample whose label is True and get its item vector (item_vec) outputted by item tower of the auxiliary model, as shown in Figure 2.
Secondly, retrieval the most relevant centroids for an item vector. The $K_{2}$ most relevant centroids for an item vector are retrieved in user memory network based on inner product. Then these selected centroids are filtered by the corresponding user profile vector to ensure positive correlations between them. If a selected centroid is negatively correlated with the user profile vector, it will be replaced by zero vector.
Thirdly, incremental update the personalized enhancement vectors. A personalized enhancement vector is predefined for each user in personalized user memory network and is initialized to zero vector. After the most $K_{2}$ relevance centroids for an item vector are gotten, the corresponding user’s personalized enhancement vector ${p_{i}}$ is updated using a small learning rate $\eta$, as shown in Formula 3.
(3) $p_{i}^{n+1}\longleftarrow arg\min_{p_{i}}\eta*\sum_{j=1}^{k_{2}}\sum_{h=1}^{d}\left(p_{ih}^{n}-\mu_{jh}\right)^{2}.$
Finally, enhancing a user’s profile with the user’s personalized enhancement vector. When estimating a user’s multiple behaviors, the personalized enhancement vector for this user is retrieved by user id in personalized user memory network and is taken as a new feature of PLE model to further enhance the user’s profile. When resources are sufficient, a personalized enhancement vector can be generated separately for each task and different media types to further improve performance.
3.4. User Consumption Sequence Enhancement (UCSE)
User history behavior sequences including valid consumption sequence (also called consumption sequence), interaction sequence, dislike sequence and so on reflect users’ recent consumption preference. As part of user interest, user history behavior sequences play an important role for the performance of ranking model. For simplicity, we take user consumption sequence as an example, as shown in Figure 1. As mentioned earlier, there is a significant difference in the number of user consumption behaviors among different users, as shown in Table 2. Therefore, to enhance users’ consumption sequence, especially for low-active users, we design a new enhancement method named UCSE, as shown in Figure 2. The basic idea of UCSE is to enhance each consumption behavior contained in user consumption sequence using the most similar one of all centroids generated by stream clustering all users’ consumed items and then produce sequence enhancement vector based on attention mechanism. The key steps of UCSE are as follows.
Firstly, getting the vector of each consumption behavior (called consumption vector for simplicity) contained in user consumption sequence. For convenience, we directly use the vector representation of consumption behavior in the main model, which include the embeddings of video ID, category 1, category 2, tag and so on. The maximum length of user consumption sequence is limited to 100. This is a short sequence that records users’ recent behaviors. Considering the average daily volume of user consumption behaviors in our scenario, this length is sufficient.
Secondly, stream clustering based on similarity between consumption vectors and centroids stored in memory network. Before training, the centroids of $M$ clusters are randomly initialized in item memory network. By calculating the similarity between consumption vector and each centroid, the consumption vector is clustered to the most similar cluster, using the similar way as Formula 1. In addition, to balance the influence of different users on clustering and reduce computation, we perform sampling based on user’s active type and about 10% of all samples take part in clustering.
Thirdly, retrieval the most similar centroid for each consumption behavior contained in consumption sequence. For predicting user’s various behaviors, the most similar centroid $\theta_{j}$ for a consumption behavior in consumption sequence is retrieved in item memory network based on similarity.
Finally, generating enhancement vector for user consumption sequence. After obtaining the most $K$ similar centroids of user consumption sequence, the enhancement vector $\upsilon_{ucse\_weighted}$ is generated based on these centroids and target item vector $\upsilon_{target}$ using Formula 4, as shown in Figure 2.
(4) $\upsilon_{ucse\_weighted}=\sum_{j=1}^{k}\frac{e^{(\theta_{j}\ *\ \upsilon_{target})}*\theta_{j}}{\sum_{j=1}^{k}e^{(\theta_{j}\ *\ \upsilon_{target})}}.$
In practice, in addition to the approach mentioned above, clustering the item vectors of positive training samples can also be used. During prediction, the user’s recent behavior sequence is used to query in the memory network that stores the cluster centers of items. Similar to Subsection 3.2, the clustering algorithm here can be replaced with other clustering algorithms, and performance can be further improved through hierarchical (tree-structured) clustering.
3.5. Extensions for Tail Items
The model performance on tail items is also poor due to highly skewed long-tail item distribution, which is usual in industrial RSs. In our recommendation scenario, among all items exposed to users within a day, the top 10% items account for over 90% of total exposures. However, the tail 90% items only account for less than 10% of total exposures. Through data analysis, we define the items whose cumulative exposure number are less than 10,000 are tail items. To improve tail-item recommendation, we employ similar methods as discussed in Subsection 3.2, 3.3 and 3.4 for tail items. The offline and online experiments conducted in Section 4 demonstrate that these approaches also yield significant improvement on tail-item recommendation. For brevity, the specific details will not be elaborated in this paper.
4. Experiments
4.1. Dataset
For offline experiments, the training and test datasets are collected from a short video recommendation scenario, which serves hundreds of millions of users. The data collected from the first 10 days is used as training data that includes about 2.9 billion samples, while the data collected from the 11th and 12th days is used as test data that consists of about 0.6 billion samples. All the compared models are trained using the training data and are tested on the test data. In addition, all models use Adam (ref34, ) as their optimizer and set batch size to 512. The values of $N$, $M$, $K1$, $K2$, $\rho$, $\eta$ are set to 256, 256, 3, 3, 0.01, 0.05 respectively. The dimensions of the vectors stored in memory networks are set to 36, 36 and 48. In addition, the dimension of personal vector is set to 4. The hyperparameters of different models are determined through offline selection. For online experiments, all the compared models are trained using the data collected during the previous two months and then are deployed in the large-scale RS to conduct A/B tests.
4.2. Evaluation Setting
4.2.1. Offline evaluation
We evaluate the performance of valid consumption task of different models using AUC (ref32, ) for comparison. For simplicity, the improvements of other tasks will not be listed. As mentioned before, if a user watches a video more than 10 seconds, the label of valid consumption is 1 (True). Otherwise, its label is 0 (False).
4.2.2. Online A/B testing
We deploy different models in our RS for one week to conduct A/B tests. We adopt User Valid Consumption (UVC) and User Duration Time (UDT) to evaluate each model, which are the two most important online metrics in our recommendation scenario. In addition, we also compare diversity metrics of the compared models, including User Consumption on Category 1 (UCC1), User Consumption on Category 2 (UCC2) and User Consumption on Tag (UCT).
UVC is the average of all users’ total valid consumptions during a day.
UDT is the average of all users’ total watching time within a day.
UCC1 is used to evaluate users’ valid consumption on category 1. UCC1 is calculated by two steps, first count the number of distinct features of category 1 of the items consumed by each user during a day, then compute the average of them.
UCC2 and UCT is calculated like UCC1 metric, which are also used to evaluate the diversity of the users’ consumption.
4.3. Compared Methods
We compare our solution with two methods (ref28, ; ref29, ) proposed by Google. Since the methods proposed by (ref3, ; ref31, ) are orthogonal to our solution, we don’t compare our solution with them. Moreover, we design four variants of our solution to illustrate the effects of each enhancement component on performance respectively.
PLE (ref18, ): PLE is proposed by Tencent for MTL which demonstrates excellent performance. In this paper, PLE is used as the benchmark for comparison.
PLE with CL: A CL enhancement component is designed for low-active users inspired by (ref28, ) based on PLE.
PLE with TL: A TL enhancement component is developed for low-active users inspired by (ref29, ) based on PLE.
PLE with UPE: UPE is implemented based on PLE, which is used for enhancing user profile for all users.
PLE with UPE and UCBE: UCBE is developed based on PLE with UPE, which is used to enhance user profile for all users from two different perspectives.
PLE with UIE: UIE including UPE, UCBE and UCSE is implemented based on PLE, which is designed to validate the performance of all enhancement components for all users.
PLE with UIE and Tail Item Enhancement (TIE): TIE is similar to user enhancement methods, as described in Section 3.5, aiming to improve tail-item recommendation.
4.4. Offline Evaluation
In this section, extensive offline experiments are conducted to demonstrate the remarkable performance of our solution compared to other methods. All the models are tested on the same dataset and the results are shown in Table 3.
Table 3. The test AUC of low-active users and high-active users of compared models. '-' represents there are no modification compared to baseline, no need for comparison.
Models Low-active Users High-active Users
PLE 0.7841 0.8023
PLE with CL 0.7845 -
PLE with TL 0.7847 -
PLE with UPE 0.7878 0.8041
PLE with UPE and UCBE 0.7897 0.8052
PLE with UIE 0.7913 0.8061
Table 4. The results of online A/B experiments. '*' represents online result of baseline. '-' represents that there are no modifications compared to baseline, no need for comparison.
Low-active Users High-active Users
Compared Models UVC UDT UCC1 UCC2 UCT UVC UDT UCC1 UCC2 UCT
PLE * * * * * * * * * *
PLE with CL +0.52% +0.37% +0.08% +0.73% +0.81% - - - - -
PLE with TL +0.79% +0.41% +0.11% +0.83% +0.90% - - - - -
PLE with UPE +3.41% +1.92% +1.13% +1.92% +2.07% +1.98% +1.34% +0.63% +1.34% +1.75%
PLE with UPE and UCBE +5.92% +2.89% +3.16% +3.20% +3.97% +3.08% +1.73% +1.22% +1.99% +2.91%
PLE with UIE +7.96% +3.65% +3.97% +4.58% +5.22% +4.06% +2.19% +1.69% +2.86% +3.27%
In offline evaluation, the AUC of PLE with CL on low-active users is slightly higher than that of PLE with an improvement of 0.04%. Due to the data augmentation approach of CL (ref28, ) easily generating meaningless data in RS, the augmentation data has limit value to the model. The AUC of PLE with TL on low-active users is 0.06% higher than that of benchmark. As previously mentioned, TL (ref29, ; ref30, ) is extremely influenced by the difference between the source and target data distributions. PLE with UPE significantly improves model performance on low-active users, with an improvement of 0.37%. Moreover, it also enhances model performance on high-active users, with an improvement of 0.18%. Through constructing personalized enhancement vectors for users from the perspective of consumption behaviors, UCBE further improves model performance on low-active users and high-active users, as shown in Table 3. In addition, UCSE enhances user consumption sequence using the most similar centroids stored in item memory network and target attention mechanism, which is particularly valuable for low-active users. Compared with other models, PLE with UIE including all enhancement components of user interest achieves the greatest improvements, increasing by 0.72% on low-active users and 0.38% on high-active users than baseline model. In addition, other tasks of PLE with UIE also show significant improvements compared with other models, which is not listed for simplicity.
It should be noted that the hyper-parameters of UIE need to be carefully selected according to specific recommendation scenario and requirements, which have a significant impact on the performance of the model. Taking PLE with UPE as an example, when different values of $N$ are set, the improvements are different, as shown in Figure 3. To balance performance and cost, $N$ is set to 256. The selection processes of other parameters are similar and will not be listed one by one.
Finally, to improve tail-item recommendation, we adopt similar approaches (TIE) to UPE and UCBE as discussed in Subsection 3.2, 3.3 and 3.4 for tail items which also significantly improves model performance on tail items, as shown in table 5.
Figure 3. The test AUC of PLE with UPE on low-active users and high-active users when N is set to different values.
Table 5. The test AUC of tail items of different models.
  Compared Models   Test AUC of Tail Items
PLE with UIE 0.7793
PLE with UIE and TIE 0.7819
4.5. Online Evaluation
For online evaluation, we deploy different models in our recommendation scenario for one week to conduct A/B tests. We take PLE as the benchmark for comparison and show the improvements of the metrics for other models. The online experiment results are shown in Table 4 and all the improvements have statistical significance with $p$-value less than 0.05.
PLE with CL and PLE with TL slightly improves UVC and UDT compared to benchmark. The reasons are detailed in Section 4.4. PLE with UPE increases +3.41% UVC, +1.92% UDT on low-active users and +1.98% UVC, +1.34% UDT on high-active users compared to benchmark, which is far superior to the above two models. Moreover, PLE with UPE also improves diversity metrics on all users, especially on low-active users, which indicates that the user’s profile is effectively supplemented. PLE with UPE and UCBE further improves online results than PLE with UPE, by constructing personalized enhancement vectors for users’ profile from the perspective of consumption behaviors. PLE with UIE achieves the best improvements compared to other models, with improvements of +7.96% UVC, +3.65% UDT on low-active users and +4.06% UVC, +2.19% UDT on high-active users. Furthermore, the diversity metrics are significantly improved by PLE with UIE as well. It is important to note that the enhancement components proposed in this paper will change the distribution of predictions outputted by ranking model. Therefore, it is recommended to adjust the weights of the fusion formula accordingly, which is also known as Multi-task Fusion (MTF) (ref35, ; ref36, ; ref37, ).
Finally, to evaluate the performance of TIE on tail items whose cumulative exposure number are less than 10,000, we compare the exposure ratio of tail items between PLE with UIE and PLE with UIE and TIE. We calculate the exposure ratio of tail items within a day for each model respectively. The exposure ratio of tail items of PLE with UIE and TIE is 1.46% higher than that of PLE with UIE, which is a significant improvement on tail items. All the modules of UIE mentioned above have been fully deployed in our RS one by one.
5. Related Work
Ranking (ref4, ; ref8, ; ref16, ; ref17, ; ref18, ; ref19, ; ref20, ; ref21, ) is responsible for predicting the scores of various user behaviors using elaborate model and lots of features, which is critical for personalized RSs. In large-scale industrial RSs such as Tencent News, YouTube, Kwai, TikTok, etc., many users only have sparse interest (also known as low-active users) because of lacking consumption behaviors including valid click, watching, like, collect, share etc. The features of user interest determine the upper bound of ranking model’s performance. Therefore, due to lack of crucial information of user interest containing user profile and user history behavior sequences, the ranking model performance on low-active users is poor. As what mentioned before, there are more than 30% of all users in our recommendation scenario whose consumption behaviors are significantly less than high-active users, resulting in their test AUC of valid consumption is nearly 2% lower than that of high-active users.
Enhancing model performance on low-active users is necessary but also quite hard. Until now, the research on low-active users is not much. (ref22, ; ref23, ; ref24, ) attempt to learn the initial ID embedding for new items/users based on meta learning. In addition, (ref25, ; ref26, ) try to get initial ID embedding for new users/items via Conditional Variational Auto-encoder (CVAE). The above methods are designed for solving cold start problem for new users/items and are not suitable for low-active users who have some consumption behaviors but not many compared with high-active users. (ref28, ) improves representation learning for sparse problem based on data augmentation and SSL framework, which can be used to enhance user interest for low-active users. However, in recommendation scenario, the data augmentation of CL can easily generate meaningless data, therefore it has limited value to the model. (ref29, ) proposes a dual transfer learning framework, which can be used to transfer knowledge from head items to tail items or high-active users to low-active users. But transfer learning is greatly influenced by the difference between the source and target data distributions and even lead to negative transfer (ref30, ). Furthermore, (ref3, ) uses personalized prior information as input and dynamically scales the bottom-level Embedding and top-level DNN hidden units, which can improve model performance on low-active users from the perspective of model structure and embedding. Based on similar idea, (ref31, ) farther optimizes ranking model performance on low-active users. However, due to lacking key interest information on low-active users, the improvements of the above two solutions are limited. To sum up, the existing works mentioned above have not achieved significant improvements for low-active users.
6. Conclusion
In this work, we first highlight the problem that lots of users in large-scale RSs merely have sparse interest, resulting in poor recommendation performance for them. To solve this difficult problem, we propose a novel user interest enhancement solution called UIE for enhancing user profile and user history behavior sequences. The key idea of UIE is to construct enhancement vectors for enhancing user interest from different perspectives, using the most similar centroids and the most relevant centroids which are generated by stream clustering and stored in different memory networks. By supplementing key information of user interest including user profile and user history behavior sequences, UIE remarkably improves the performance of ranking model on low-active users and achieves significant improvements on high-active users as well. For offline evaluation, we conduct extensive experiments to compare the performance of different models. The results show that UIE significantly exceeds other methods, especially on low-active users. Furthermore, we also conduct online A/B testing in a large-scale RS and the results demonstrate our model performs significantly better than other models, resulting in +7.96% UVC, +3.65% UDT on low-active users and +4.06% UVC, +2.19% UDT on high-active users. In addition, our model also improves diversity metrics on low-active and high-active users significantly. Similar methods are used for tail items as well to improve their exposure rate and recommendation performance. UIE has been deployed in multiple large-scale RSs and has inspired subsequent relevant works.

### v6

@@
 Furthermore, we extend our approach to long-tail items using similar methods, which also yields excellent improvements.
-We conduct extensive offline and online experiments in a large-scale industrial RS.
-The results demonstrate that our model substantially outperforms other existing approaches, especially for users with sparse interest.
+We conduct extensive offline and online experiments in an industrial RS.
+The results demonstrate that UIE substantially outperforms other existing approaches, especially for users with sparse interest.
 UIE has been deployed in several large-scale RSs at Tencent since 2022, which was made public on 21 May 2024.
 In addition, UIE-based methods have also been successfully applied in candidate generation, pre-ranking, and context-DNN stages.
-Multiple teams have developed solutions based on UIE, focusing primarily on updating clustering algorithms and attention mechanisms.
-As far as we know, UIE has been deployed by many companies.
+Multiple teams have developed solutions based on UIE, focusing on updating clustering algorithms and attention mechanisms.
+As far as we know, UIE has been applied in multiple RSs, advertising systems and search engines.
 The thoughts of UIE, dynamic streaming clustering and similarity enhancement, have inspired subsequent relevant works.
 Recommender Systems, User Interest Enhancement, Stream Clustering and Memory Networks, Long-Tail Optimization, Multi-Task Learning
-journalyear: 2024copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; May 21, 2024; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’24), May 21, 2024, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/24/05doi: 10.1145/3690624.3709429ccs: Information systems Recommender systemsccs: Information systems Personalization
+journalyear: 2024copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; May 21, 2024; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’24), May 21, 2024, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/24/05doi: 10.48550/arXiv.2405.13238ccs: Information systems Recommender systemsccs: Information systems Personalization
 1. Introduction
@@
 The weighted vector and personal vector are concatenated and used as the input of the generalization layer, as shown in Figure 2.
+We adopt several structures similar to PEPNet (ref3, ) as the generalization layer to further improve performance.
+The generalization layers of the following components also adopt the similar approach.
 (2) $\upsilon_{weighted}=\sum_{j=1}^{k_{1}}\frac{e^{(\mu_{j}\ *\ \upsilon)}*\mu_{j}}{\sum_{j=1}^{k_{1}}e^{(\mu_{j}\ *\ \upsilon)}}.$
@@
 UIE has been deployed in multiple large-scale RSs and has inspired subsequent relevant works.
+Furthermore, the thoughts of UIE have also been applied in advertising systems and search engines.